### v2

# MobilityGPT: Enhanced Human Mobility Modeling with a GPT model

Abstract
Generative models have shown promising results in capturing human mobility characteristics and generating synthetic trajectories. However, it remains challenging to ensure that the generated geospatial mobility data is semantically realistic, including consistent location sequences, and reflects real-world characteristics, such as constraining on geospatial limits. We reformat human mobility modeling as an autoregressive generation task to address these issues, leveraging the Generative Pre-trained Transformer (GPT) architecture. To ensure its controllable generation to alleviate the above challenges, we propose a geospatially-aware generative model, MobilityGPT. We propose a gravity-based sampling method to train a transformer for semantic sequence similarity. Then, we constrained the training process via a road connectivity matrix that provides the connectivity of sequences in trajectory generation, thereby keeping generated trajectories in geospatial limits. Lastly, we proposed to construct a preference dataset for fine-tuning MobilityGPT via Reinforcement Learning from Trajectory Feedback (RLTF) mechanism, which minimizes the travel distance between training and the synthetically generated trajectories. Experiments on real-world datasets demonstrate MobilityGPT’s superior performance over state-of-the-art methods in generating high-quality mobility trajectories that are closest to real data in terms of origin-destination similarity, trip length, travel radius, link, and gravity distributions.
1 Introduction
The widespread integration of location-based services and smart GPS devices, such as smartphones and watches, has made continuous monitoring of human mobility both desirable and feasible. These technologies capture diverse and detailed human movement information, with mobility trajectories representing the finest granularity of individual-level mobility characteristics. Such trajectories are crucial in various applications, including mobility modeling, commercial business analysis, and disease spread control (Siła-Nowicka et al., 2016; Changruenngam et al., 2020).
Despite increasing demand for human mobility trajectory datasets, numerous challenges hinder their access and distribution (Kong et al., 2023). First, these datasets are typically collected by private companies or government agencies, hence presenting privacy concerns due to the potential disclosure of an individual’s sensitive lifestyle patterns (e.g., home or office addresses and points of interest). Legal implications, such as those outlined in the California Data Privacy Act and GDPR, highlight the importance of carefully handling spatiotemporal traces. Second, datasets owned by companies may expose proprietary business models and are often inaccessible for research purposes (Wang et al., 2019). Lastly, publicly available datasets often lack diversity or quality, with gaps in data points and intrinsic noise that significantly reduce their utility. These limitations impede the progress of urban planning or transportation research (Luca et al., 2021). Therefore, it is imperative to establish alternative trajectory data sources that are both high-quality and accessible for research purposes.
Figure 1: Generative models that do not incorporate mobility characteristics struggle to capture human mobility patterns accurately.
Current generative models designed for human mobility exhibit multiple drawbacks: they fail to capture the sequential mobility characteristics, often lack the continuity of generated trajectories, and do not consistently adhere to geospatial constraints. Approaches based on GANs (Goodfellow et al., 2020) and VAEs (Kingma & Welling, 2013) structure data in a tabular format, which, unfortunately, fails to preserve the inherent correlations among locations. Models like LSTMs, and sequential GANs (Yu et al., 2017) struggle to maintain the smooth continuity of locations and capture realistic human mobility characteristics. A further limitation of these methods lies in the elevated mismatches observed in geospatial location sequences. In short, a successful generative model for trajectory datasets should precisely grasp the intrinsic spatial-temporal behaviors.
Large language models (LLM), such as GPT (Generative Pre-trained Transformer), offer a robust foundation for sequence generation tasks and can be refined through fine-tuning to cater to distinct objectives. Trajectories and sentences share several similarities that make GPTs a promising approach for trajectory generation. Both consist of ordered sets of elements chosen from finite pools (road links and words, respectively). They exhibit semantic or spatio-temporal relationships, adhering to distinct rule systems such as language rules for sentences and geographical constraints for trajectories. Due to these parallels, the techniques developed for natural language processing can be adapted to model and generate realistic trajectories. (Musleh et al., 2022) Besides, the gravity model holds significance in human mobility modeling as it provides a structured framework for estimating and understanding the flow of interactions between different locations (Zipf, 1946). Using gravity as part of human mobility modeling could bring further advantages to synthetic trajectory generation tasks.
In this study, we leverage the power of GPT for synthetic trajectory generation. However, directly training GPT on mobility data without geospatial insights would generate unrealistic sequences. To generate more realistic synthetic trajectories, we introduced several innovative methods, including (1) gravity-aware sampling, which incorporates gravity modeling of trajectory data into training updates, and (2) the use of road connectivity matrix (RCM) masking to eliminate disconnected location sequences from logits (see Figure 1). Pretraining GPT models on trajectory data captures the general sequence characteristics in an unstructured manner. We, furthermore, proposed an automated fine-tuning pipeline that improves the trajectory quality by leveraging transportation-specific metrics to evaluate and optimize the generated sequences without relying on human labeling. Our multi-objective framework, MobilityGPT, not only extends the application of GPT models to human mobility modeling but also collectively enhances the model’s capacity to capture intricate patterns and adhere to realistic geospatial constraints (e.g., trajectories consist of road segments that are indeed connected). Our contributions can be summarized as follows:
  • We propose an LLM-inspired MobilityGPT human mobility modeling approach incorporating multiple unique features with a geospatially aware pretraining method and automated fine-tuning approach without human feedback.
  • While gravity-aware sampling trains a generative model with respect to Origin-Destination (OD) pair-gravity values, the next sequence prediction is conditioned on the RCM that preserves the continuity of the generated trajectories.
  • We propose a novel method for constructing a preference dataset to fine-tune MobilityGPT using reinforcement learning, enhancing the similarity of generated trajectories in terms of travel length.
  • Through comprehensive experimental analysis using real datasets, we demonstrated that the proposed methods can generate trajectories with high fidelity while preserving essential statistical and semantic properties of human mobility.
2 Related Work
Generating synthetic mobility trajectories from real data is a promising approach for protecting sensitive information (Kapp et al., 2023). Traditional model-based methods, such as those parameterizing human mobility, have limitations due to the complexity of mobility characteristics. Early data-driven approaches, employing techniques like gravity models and decision trees (Pappalardo et al., 2016; Jiang et al., 2016), face challenges in capturing sequential transitions between locations. In contrast, model-free techniques, particularly generative machine learning models, have gained traction for their ability to learn from data without external input parameters (Baratchi et al., 2014). Such generative models have been explored for human mobility modeling using various data representations, including grids in tabular format (Ouyang et al., 2018), image-like trajectory modeling (Cao & Li, 2021), and sequential grid format (Feng et al., 2020). However, these models exhibit limitations in generating trajectories that accurately capture geospatial complexity.
One promising generative model direction relies on GPS sampling. Due to mismatches, the LSTM-based trajectory generation model cannot provide continuous motion of human mobility in diverse geospatial areas (Rao et al., 2020). A similar continuous GPS trajectory generation method with a U-Net neural network using Diffusion models is proposed in (Zhu et al., 2024). Another generative direction with ML for mobility trajectories is employing two-stage training methods (Wang et al., 2021; Jiang et al., 2023; Long et al., 2023). A recent GAN-based model, relying on road links with the $A^{*}$ path-finding algorithm between regions employing two-stage GANs, has shown promise in identifying diverse paths for specified origin-destination pairs (Jiang et al., 2023). However, it falls short in generating synthetic trajectories, as the generation process initiates with an origin-destination pair from testing trajectories. Finally, sequential GAN methods have also been employed for mobility modeling with different variants (Yu et al., 2017; Feng et al., 2020). Unlike previous methods, MobilityGPT leverages the auto-regressive GPT model for trajectory generation, enabling a comprehensive exploration of spatial and temporal distributions with high-quality outputs by incorporating road links into sequence modeling.
The transformer-based GPT concept has been applied to diverse applications, from vision (Chen et al., 2020), music (Banar & Colton, 2022) to network data (Kholgh & Kostakos, 2023). In addition, researchers have made great progress on various fine-tuning models for enhancing the GPT model capabilities (Zhang et al., 2019; Brown et al., 2020; Lambert et al., 2022). However, these fine-tuning models mainly rely on human feedback that evaluates the quality of the generated sequences. This study proposes a geospatially-aware GPT model, MobilityGPT, with an innovative fine-tuning approach for generating synthetic mobility trajectories.
3 Preliminaries
3.1 Discretization with Map-Matching
Let $D(V,E)$ represent the road network as a weighted digraph, where the set of nodes $V$ correspond to road intersection, set of edges $E$ to roads, and weights representing link metrics, such as length of the link or traffic volume. There are two sets of trajectories: GPS trajectories and link trajectories.
1) GPS trajectories: A sequence of GPS coordinates with $l$ number of samples $\mathbf{x}\in\mathcal{G}=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{l}\}$ forms a GPS trajectory that reflects the continuous motion of the object.
2) Link trajectories: Given the $s$ number of vehicles on the road network, each vehicle travels between origins and destinations using an ordered link path generating a user travel path known as a micro-graph $\Phi\in D$. Every link trajectory has $n$ number of links $\phi\in\Phi=\{\phi_{1},\phi_{2},...,\phi_{n}\}$ and $\Phi\subset{E}$. The set of trajectories is the corpus of all link trajectories with $s$ users $\Phi\in\Lambda=\{\Phi_{1},\Phi_{2},...,\Phi_{s}\}$.
The map-matching process is crucial in refining GPS coordinates, which provide an approximate device location but may not consistently represent the exact travel path due to various errors. Map-matching generates an ordered set of road network links describing the user’s trajectory after accounting for the underlying road network $D(V,E)$ and GPS points (Quddus et al., 2007). Addressing location sequences in the continuous domain is challenging in conventional human mobility modeling, often resorting to a discretized grid domain. However, this grid-like representation does not entirely eliminate geospatial mismatching risks. To overcome this, our MobilityGPT model discretizes locations into road segments with map-matching, resulting in more meaningful and realistic generated trajectories.
3.2 Problem Definition
In an urban setting, a generative model is tasked with learning the intrinsic mobility characteristics from a provided set of link trajectories, $\Phi\in\Lambda=\{\Phi_{1},\Phi_{2},...,\Phi_{s}\}$. The objective is to generate synthetic trajectories that closely resemble the training trajectories in terms of various utility metrics, thereby capturing and reproducing the key features of mobility present in the input data.
4 Mobility Modeling with Generative Transformers
This section describes the components of the proposed MobilityGPT for generating synthetic high-fidelity trajectories. MobilityGPT achieves this through a multi-objective training pipeline. Figure 2 illustrates the framework of our MobilityGPT mechanism, which can be divided into two parts: pretraining for learning sequence characteristics and fine-tuning for enhancing the quality of generated trajectories. The pretraining stage (Step 1 in Figure 2) learns the mobility modeling on the road-link sequences using the gravity of training data and RCM. The fine-tuning stage (Steps 2 and 3 in Figure 2) aims to improve the similarity of synthetic trajectories to the training trajectory samples in terms of travel length.
Figure 2: Overview. The proposed MobilityGPT framework employs Generative Pretrained Transformer (GPT) architectures and self-adaptive reinforcement learning (RL) fine-tuning methods within a human mobility-aware training pipeline for synthetic trajectory generation. The process involves three steps: 1) Pretraining the GPT model with a gravity model and road connectivity matrix, 2) Constructing a trajectory-driven reward model, and 3) Fine-tuning the MobilityGPT pre-trained model using RL policy optimization methods with the trained reward model providing valuable feedback.
4.1 Generative Transformers
Mobility trajectories refer to the sequential records of locations or movements over time, capturing the spatiotemporal patterns of an entity’s mobility. Their characteristics play a crucial role in human mobility modeling, providing insights into individual or collective movement behaviors.
Recent breakthroughs in NLP have proven that transformer architectures are remarkably effective in processing word sequences (Zhao et al., 2023). We have found that trajectory sequences, representing the spatiotemporal movements of individuals or objects, share similarities with sentences across four key dimensions: sequential dependencies, spatial relationships, contextual embeddings, and variable-length sequences. First, trajectory sequences exhibit sequential dependencies similar to sentences, where the order of locations matters. Additionally, trajectories involve spatial relationships between locations, just as words in a sentence convey semantic relationships. Furthermore, trajectory sequences, like sentences, can have variable lengths. Finally, trajectories benefit from contextual embeddings that consider the entire sequence, similar to the contextual understanding of words in sentences.
Given these strong similarities between trajectories and word sequences, MobilityGPT benefits two key elements of transformer architecture: self-attention and autoregressive generation. It allows efficient handling of sequential trajectory data by capturing long-range dependencies and generating contextually coherent sequences.
The self-attention mechanism computes a weighted sum of values $V$ based on attention scores $A$ assigned to each element in the input sequence:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)\cdot V$ (1)
where $Q$, $K$, and $V$ represent the query, key, and value matrices. The softmax operation normalizes the attention scores, and $d_{k}$ is the dimensionality of the key vectors. The weighted sum is then used as the output of the self-attention mechanism.
In an autoregressive nature of transformers, the probability of predicting the next element $y_{t}$ in the sequence given the context $Y_{<t}$ is computed using the chain rule of probability:
$P(y_{t}|Y_{<t})=P(y_{1})\cdot P(y_{2}|y_{1})\cdot\ldots\cdot P(y_{t}|y_{t-1})$ (2)
By formulating human mobility trajectories as a sequence generation task with road links from $D(V,E)$, transformer-based models can be applied to generate realistic and contextually relevant synthetic trajectories given the real trajectories. The tokenization of road links, combined with the self-attention mechanism, enables the model to capture spatial dependencies and generate coherent and meaningful sequences of road links.
4.2 Tokenizer Modeling
In MobilityGPT, tokens symbolize locations in terms of road links, and the contextual embeddings derived from tokenization contribute to generating embeddings that consider the holistic context of the entire trajectory. For each trajectory, we incorporate an “end of trajectory” token, denoted as <EOT>, which is appended to each trajectory during training. Given a trajectory sequence $\phi\in\Phi=\{\phi_{1},\phi_{2},...,\phi_{n}\}$ in terms of road links, where $\phi_{i}$ denotes the $i$-th token in the trajectory. The tokenization process is defined as follows:
[…]
This inclusion of <EOT> serves a dual purpose. First, it acts as a sentinel token to signify the end of a trajectory. This is expressed as if $\text{if }\phi_{i}=\texttt{<EOT>}\text{, then trajectory }i+1\text{ begins}$. The model leverages this information to capture the distinct OD points of each trajectory effectively. Second, this tokenization allows for generating diverse and randomized trajectory sequences. A detailed discussion can be found in Appendix C.3.
Utilizing this tokenization strategy is integral to our model’s ability to capture spatial patterns, differentiate between trajectories, and generate realistic sequences with varied ODs. This formulation contributes to the robustness and versatility of our trajectory generation approach.
4.3 Map Understanding with Gravity Model and Road Connectivity Matrix
Transformer-based models capture sequential information within long data sequences. However, human mobility modeling presents distinct challenges, necessitating a comprehensive understanding of geographical attributes and constraints. For instance, the sequence of links does not contain hidden attributes of human mobility, such as regional mobility flows (traveling from eastbound of city to west), traffic rules (must right-turn, no U-turn), and trip length. We have developed two innovative methods to overcome these challenges: the Gravity Model and the Road Connectivity Matrix (RCM). The Gravity Model enhances the transformer’s capability to understand the mobility flow patterns between regions, providing a context of spatial importance. Together, RCM allows the transformer to learn the complex structure of the road network $D(V,E)$. These strategies collectively train the transformer with an understanding of geospatial information without needing auxiliary data.
1) Gravity model.The area is discretized into regions $r\in\mathcal{R}$, with $\text{RegionW}(r_{x})$ calculated as the number of trajectories starting or ending in the region $r_{x}$. The gravity value $\text{Gravity}(r_{x},r_{y})$ quantifies traffic flow between origin $r_{x}$ and destination $r_{y}$ regions as:
$\text{Gravity}(r_{x},r_{y})=\frac{\text{RegionW}(r_{x})\times\text{RegionW}(r_{y})}{d^{2}(r_{x},r_{y})}$
where $d^{2}(r_{x},r_{y})$ is the squared Euclidean distance between regions. Gravity is based on the principle that the interaction between two regions is directly proportional to the product of their respective weights and inversely proportional to the square of the distance between them. During training, as shown in Figure 2, a gravity-aware sampler selects trajectory sequences $\Phi=\{\phi_{1},\phi_{2},...,\phi_{n}\}$ based on $\text{Gravity}(r_{x},r_{y})$, with $\phi_{1}\in r_{x}$ and $\phi_{n}\in r_{y}$.
2) Road connectivity matrix. The clarity of road connectivity is often diminished when using GPS trajectories compared to traditional 2D maps. This is because, in some instances, roads that appear connected on a map might not be so in reality due to physical barriers or design differences, such as one-way streets and pedestrian zones. To mitigate this issue, we propose a method to precisely extract road connectivities from real-world data and incorporate this information into the training process of MobilityGPT. To achieve this, we define the Road Connectivity Matrix (RCM) based on the links from link trajectories, denoted as $\Phi$:
$RC_{\phi_{x},\phi_{y}}=\begin{cases}1&\text{if }(\phi_{x},\phi_{y})\text{ is aconsecutive pair in }\Phi,\\ &\quad\text{or}\\ &\text{if }\phi_{x}\text{ or }\phi_{y}\text{ is in the last row/column,}\\ 0&\text{otherwise}.\end{cases}$ (4)
RCM effectively represents the connectivity between different road links. If two links, $\phi_{x}$ and $\phi_{y}$, are directly connected in the real world, their corresponding element in RCM is set to 1; otherwise, it is set to 0. Additionally, one row and one column, both filled with 1s, are added to the matrix to facilitate the generation of <EOT>. Specifically, the last linear layer of the transformer, denoted as $F$, is adjusted by the RCM prior to the softmax activation function:
$F^{\prime}=F\cdot RC$ (5)
$\text{Softmax}(F^{\prime})=\frac{e^{F^{\prime}}}{\sum e^{F^{\prime}}}$ (6)
this integration is visualized in Figure 2, showing how the RCM modifies the output of the transformer’s last linear layer. By multiplying $F$ with the RCM, we ensure that only the connections between actually connected roads are considered. This process filters out any inaccuracies in road connectivity that might have been introduced by GPS data, thus enabling the model to have a more accurate understanding of the actual road network.
4.4 Reinforcement learning from trajectory feedback (RLTF)
Fine-tuning is essential for sequence generative models (Zhang et al., 2023). Pre-trained transformer architectures provide a foundation, but fine-tuning customizes models to specific objectives, such as similarity of trajectories, enabling them to learn complex patterns and nuances. This process is particularly vital in applications like human mobility, where capturing intricate movement patterns is crucial. Preference datasets are crucial for fine-tuning, allowing domain knowledge and critique to be integrated into training. Building on this, we propose a novel approach to construct a trajectory-aware preference dataset that combines established fine-tuning strategies with a new preference feedback dataset - all without requiring human labeling. This enables the model to self-improve based on its output feedback, minimizing human intervention and addressing scalability and bias challenges (Ouyang et al., 2022).
Figure 3: The scheme of fine-tuning dataset generation for Reinforcement Learning from Trajectory Feedback.
Reinforcement learning has proven to be a powerful approach for capturing the long-term dependencies and spatiotemporal characteristics inherent in human mobility modeling, owing to the complexity of this task (Reed et al., 2022). The proposed approach, incorporating our novel preference dataset with reinforcement learning, termed Reinforcement Learning from Trajectory Feedback (RLTF), in an automated feedback process, involves several key steps. First, a reward dataset is created by pairing prompts (partial trajectories) with compilations (complete trajectories generated by the model) and assigning scores based on the similarity of the compilation lengths to the original trajectory lengths (see Figure 3). This reward dataset is then utilized to train a reward model $\mathbf{U}$ through supervised learning. Subsequently, the trained reward model $\mathbf{U}$ is employed to fine-tune the MobilityGPT model through a Proximal Policy Optimization (PPO) policy $\pi$, effectively leveraging reinforcement learning without requiring human feedback.
Preference dataset construction for RLTF. Instead of relying on human feedback in fine-tuning of MobilityGPT, we propose using quantitative metrics to construct the preference dataset. Given a partial trajectory prompt, we generate two complete trajectories and assess their similarity to the actual trajectory by assigning a trajectory similarity score based on the relative lengths of the generated and reference trajectories from the training set. This score measures how well the generated trajectories match the characteristics, particularly length, of the training data, serving as a proxy for human preference without requiring human labeling.
The process, as described in Figure 3, begins with a full trajectory $\Phi={\phi_{1},\phi_{2},\ldots,\phi_{n}}$ of length $L(\Phi)$. The first $m$ road-links $\Phi^{\prime}_{0}={\phi_{1},\phi_{2},\ldots,\phi_{m}}$ serve as a prompt for MobilityGPT to predict two possible continuations: $\Phi^{\prime}_{1}={\phi_{m+1},\ldots,\phi_{k}}$ and $\Phi^{\prime}_{2}={\phi_{m+1},\ldots,\phi_{z}}$. These are concatenated with $\Phi^{\prime}_{0}$ to form two complete trajectories, evaluated using $\Gamma_{1}=|L(\Phi)-L(\Phi^{\prime}_{0}+\Phi^{\prime}_{2})|$ and $\Gamma_{2}=|L(\Phi)-L(\Phi^{\prime}0+\Phi^{\prime}1)|$. The trajectory with lower $\Gamma$ is labeled $\Phi_{\text{chosen}}$, and the other $\Phi_{\text{rejected}}$, forming a preference dataset for RLTF. By leveraging this proximity-based metric, we eliminate the need for human labeling, streamlining the dataset creation process and potentially mitigating biases introduced by human subjectivity. The details of reward model training and RL-based fine-tuning methods are explained in Appendix A with more details. It is also worth noting that MobilityGPT can be generalized to different fine-tuning strategies with the proposed preference dataset.
5 Experiments
MobilityGPT’s effectiveness is validated through extensive experiments on real-world datasets against recent benchmarks. We will release our codes and data upon acceptance.
5.1 Experimental setup
1) Dataset description. We have used two datasets of GPS trajectories from taxi drivers in Porto, Portugal, and Beijing, China. The Porto dataset has a sampling period of $15$ seconds, whereas the Beijing dataset has a sampling period of $1$ minute. Both datasets have been preprocessed and map-matched by (Jiang et al., 2023) to remove any errors. The Porto dataset contains $695,085$ unique trips on $11,095$ road links, and the Beijing dataset contains $956,070$ trajectories on $40,306$ road links.
2) Benchmark models. We assess the performance of MobilityGPT with respect to four synthetic trajectory generation models as baselines. We carefully selected the benchmarks as one from each distinct approach: discrete sequential GAN (SeqGAN (Yu et al., 2017)), continuous GPS generation (LSTM-TrajGAN (Rao et al., 2020)), two stage generative model (TS-TrajGen (Jiang et al., 2023)) and continuous diffusion generative model (DiffTraj (Zhu et al., 2024)). The details of these benchmark models can be found in Appendix B.1.
Table 1: Comparison of the utility metrics with benchmark studies. Bold and underlined results show the best and second-best results, respectively. For all metrics, lower values indicate better performance. While query error relies on normalized error rates for most visited places, other metrics use the Jensen-Shannon distributional divergence statistic.
  Porto-Taxi BJ-Taxi
Methods Query E. OD Trip L. Radius Gravity Query E. OD Trip L. Radius Gravity
SeqGAN 0.138 0.116 0.220 0.191 0.326 0.063 0.296 0.162 0.275 0.120
LSTM-TrajGAN 0.234 0.254 0.435 0.425 0.277 0.078 0.182 0.398 0.365 0.219
TS-TrajGen 0.309 0.120 0.161 0.115 0.248 0.277 0.105 0.570 0.485 0.336
DiffTraj 0.267 0.177 0.126 0.123 0.252 0.074 0.146 0.158 0.123 0.114
MobilityGPT 0.144 0.114 0.124 0.107 0.225 0.066 0.099 0.123 0.102 0.105
5.2 Evaluation metrics
Our first utility metric, mainly used for graph data, is query error, which quantifies the mobility characteristics of frequently visited places (Chen et al., 2011). A lower query error indicates that the output synthetic trajectories are more similar to the training trajectories.
The Jensen-Shannon divergence (JSD) is a widely recognized similarity metric primarily employed for quantifying the similarity between two probability distributions (Lin, 1991). Additionally, we evaluate our proposed MobilityGPT using JSD for four metrics: OD similarity, trip length, travel radius, and gravity distribution. Please refer to Appendix B.2 for a more detailed description of our evaluation metrics.
In addition to divergence metrics, we introduce a new metric that measures link trajectory connectivity as a percentage. We employed the connectivity metric in our ablation studies to assess the impact of various components on link trajectory connectivity, as shown in Table 5.4a.
5.3 Comparison with benchmark generative methods
Table 1 summarizes the performance of MobilityGPT and benchmark models. We sampled $5000$ random trajectories from the test set and generated same number of synthetic trajectories. The results show the significant advantage of MobilityGPT over existing state-of-the-art models for trajectory generation. Across both Porto-Taxi and BJ-Taxi, MobilityGPT consistently outperforms all baseline models, achieving the best or second-best performance on multiple utility metrics.
Notably, MobilityGPT demonstrates up to $24.07\%$ improvement over the second-best model in modeling trip lengths and up to $17.07\%$ improvement in capturing spatial radius on BJ-Taxi. Additionally, it exhibits up to $9.27\%$ improvement in capturing gravity patterns on Porto-Taxi. MobilityGPT excels in accurately modeling intricate characteristics of individual trajectories, such as spatial and temporal patterns, by effectively capturing the relationships between locations and trajectories, addressing a significant limitation of previous approaches. Its remarkable performance in modeling OD distributions highlights its capability to capture underlying patterns governing traffic flow and human mobility. Overall, MobilityGPT’s superior performance across multiple metrics highlights its effectiveness as a comprehensive and powerful trajectory generation model, positioning it as a valuable tool for applications in transportation, urban planning, and mobility analysis.
The primary objective of mobility generative models is to produce trajectories that closely resemble the original data by capturing the overall mobility characteristics, rather than generating exact replicas. Figure 4 depicts the spatial densities generated by baseline methods for Porto City. While all the maps demonstrate that the generated trajectories accurately capture the city’s geographic profiles, MobilityGPT has a clear representation of geospatial density in the generated trajectories compared to the original test trajectories. Clearly, MobilityGPT produces trajectories that closely resemble the originals both in urban areas and the outskirts of the city, surpassing other methods significantly.
5.4 Ablation study
As MobilityGPT incorporates three distinct contributions to the standard GPT model for comprehensively capturing human mobility patterns – the gravity model, RCM, and RL fine-tuning (RLTF) – we performed experiments using different variants of these models, evaluated using three comparison metrics as shown in Table 5.4a. On average, with the additional components of MobilityGPT, we observe a $7.4\%$ improvement in trajectory similarities in terms of ODs and a $19.5\%$ improvement over trip length similarities. Furthermore, without the RCM, MobilityGPT fails to preserve connectivity in the generated trajectories. RCM ensures connectivity by conditioning the generation of the next link on connected links, effectively capturing the spatial constraints inherent in human mobility patterns.
We compare different fine-tuning methods with our proposed preference dataset construction approach: supervised fine-tuning (SFT) and direct preference optimization (DPO) (Rafailov et al., 2024) and summarized their performance in Table 5.4b. We found that RLTF is effective for capturing the complexity of mobility modeling and outperforms other fine-tuning strategies.
Table 2: Ablation study shows the influence of (a) removing each component on the performance of MobilityGPT and (b) on different methods for fine-tuning.
       RCM Gravity Model RLTF OD Trip L. Conn.
0.123 0.154 0.87
0.122 0.146 0.83
0.120 0.147 1.0
0.114 0.140 1.0
0.114 0.124 1.0
(a)        SFT DPO RLTF OD Trip Length 0.116 0.171 0.115 0.163 0.114 0.124   (b)
(a) Original Trajectories
(b) MobilityGPT
(c) TS-TrajGen
(d) SeqGAN
(e) LSTM-TrajGAN
(f) DiffTraj
Figure 4: Visual representation of the geospatial density of different methods. MobilityGPT exhibits the highest similarity to the original trajectories.
6 Discussion
This paper presents a GPT-based generative method for modeling human mobility characteristics. The proposed mechanism considers a multi-objective learning method with pre-training and fine-tuning stages. MobilityGPT pre-training leverages the understanding of human mobility using the string self-attention structures by conditioning the trajectory sampling with gravity model and sequence sampling with road connectivity matrix. To further enhance the MobilityGPT capabilities for generating similar trajectories to real trajectories, we propose a fine-tuning strategy that does not require a human assessment. The novel fine-tuning method trains a reward using the trajectory trip length and employs the reward model as an import to the RL policy gradient method. The experiments validate the effectiveness of the proposed MobilityGPT approach for generating realistic mobility trajectories from real trajectory samples by comparing with four state-of-the-art models.
Limitation: MobilityGPT achieves state-of-the-art results while handling road network constraints, but lacks formal privacy guarantees for synthetic trajectories. Future directions involve improving generalizability and incorporating practical privacy assurances.
Appendix A Reinforcement Learning from Trajectory Feedback Details
This section provides details about the novel RLTF fine-tuning strategy for generating high-quality synthetic mobility trajectories.
A.1 Reward model training.
A Reward (preference) Model $\mathbf{U}$ is trained using the proposed preference dataset to predict the probability that a given trajectory is chosen over a rejected alternative. The loss function is formulated as follows:
$-\log(\sigma(\mathbf{U}(\Phi_{\text{chosen}})-\mathbf{U}(\Phi_{\text{rejected}})))$ (7)
The goal of the loss function is to maximize the probability that the chosen trajectory has a higher score, thereby encouraging the model to assign a larger difference between $\Phi_{\text{chosen}}$ and $\Phi_{\text{rejected}}$. As the training progresses, the model becomes more capable of recognizing the characteristics of trajectories that align with the preferred trip length.
A.2 Fine-tuning with reinforcement learning.
Given the pre-trained MobilityGPT model and supervised trained reward model, we train an RL policy that optimizes the MobilityGPT model for generating higher-quality synthetic trajectories. Since policy learning relies on the reward that policy receives from the environment, the reward model provides feedback for RL policy optimization. We followed a general policy optimization method as in (Ziegler et al., 2019). The RL agent learns a fine-tuning policy $\pi$, which is optimized with the well-known Proximal Policy Optimization (PPO) method (Schulman et al., 2017). The reward of policy $\pi$ consists of two terms: reward model logits given the prompt $\Phi^{\prime}_{p}$ and compilation $\Phi^{\prime}_{c}$ pair received from the MobilityGPT as $\mathbf{U}(\Phi^{\prime}_{p},\Phi^{\prime}_{c})$ and policy shift constraint in terms of the KL divergence given the baseline policy $\pi$. The full reward can be represented as follows:
$\text{R}(\Phi^{\prime}_{p},\Phi^{\prime}_{c})=\mathbf{U}(\Phi^{\prime}_{p},\Phi^{\prime}_{c})-\beta\log\left(\frac{\pi^{\text{PPO}}_{\theta^{\prime}}(\Phi^{\prime}_{c}|\Phi^{\prime}_{p})}{\pi^{\text{base}}(\Phi^{\prime}_{c}|\Phi^{\prime}_{p})}\right).$ (8)
It is worth emphasizing that the neural network architecture of the reward model $\mathbf{U}$ can be simplified to enhance efficiency and interpretability. Moreover, there is the option to incorporate an online fine-tuning stage, where the policy $\pi$ and the reward model $\mathbf{U}$ are periodically re-trained. However, in our experimental assessments, we did not delve into the performance of the retraining process. This aspect and the formal exploration of privacy-preserving ML models are left for future research endeavors.
Appendix B Experimental Setup Details
This section describes the benchmark models and evaluation metrics in detail.
B.1 Benchmark models
We performed experiments on original implementation hyperparameters for all benchmark models using their publicly available source codes. We experimented with similar structured generative models with our proposed MobilityGPT model, including discrete and continuous structure formats.
  • SeqGAN (Yu et al., 2017): SeqGAN is a state-of-the-art sequential generative model that trains trajectories in discrete sequences with RL policy gradient methods using Monte Carlo sampling.
  • LSTM-TrajGAN (Rao et al., 2020): This is a tabular LSTM model by learning the normalized GPS location divergences with respect to the center of the dataset.
  • TS-TrajGen (Jiang et al., 2023): TS-TrajGen is a method based on GANs that incorporates a modified $A*$ path search algorithm. This adjustment ensures spatial continuity in the generated data. Additionally, the method includes topological constraints to align the generated trajectories with existing road networks, ensuring road matching. TS-TrajGen requires fixed OD pairs as inputs. We use OD pairs from the real training set as inputs during inference.
  • DiffTraj (Zhu et al., 2024): is the most recent and promising generative model for mobility trajectories. DiffTraj employs a diffusion model in the U-Net structure for generating GPS trajectories.
B.2 Evalaution metrics
To evaluate the performance of the generative models, we need a proper comparison metric with a wide range of mobility characteristics that quantifies the quality of generated trajectories given the real trajectories. In this project, we employed four well-known comparison metrics from literature (Hasan et al., 2013; Gursoy et al., 2018) in addition to our proposed gravity and connectivity utility metrics. These metrics have been widely used for the utility of synthetic mobility data generation models (Haydari et al., 2022; Zhu et al., 2024).
Query Error: The query error, mainly used for evaluating data synthesis algorithms, is a popular metric for synthetic data generation models. We use spatial counting queries in the form of “the number of trajectories passing through a certain road-link”. Given road network $D(V,E)$, we uniformly sampled $500$ road links from the road network. Then, the normalized absolute difference between the number of real and synthetic trajectories passing through each of these links is computed by the following:
$\hbox{QE}(f(\Sigma))=\frac{|f(\Psi)-f(\Sigma)|}{\max\left\{f(\Psi),s\right\}},$ (9)
where $f(\Psi)$ and $f(\Sigma)$ stand for the query outcome from the original and synthetic trajectories, respectively, and $s$ is sanity bound for mitigating the effect of the extremely small selective queries. We specified the sanity bound $s$ as $1\%$ of the users.
Apart from the query error, the main utility metrics inspect the distributional similarities for synthetic trajectories with respect to real trajectories that train the generative model. Jensen-Shannon divergence (JSD) is a well-known similarity metric mainly used for measuring the similarity of two probability distributions (Lin, 1991). We employ JSD for several human mobility characteristics.
We can look at overall distributions of the generated trajectories $\Lambda_{gen}$ and real trajectories $\Lambda_{real}$, and JSD can provide a summary statistic for a pair of distributions. These are the main distributions we extract from the $\Lambda_{gen}$ and $\Lambda_{real}$:
  • OD: This metric evaluates the origin-destination similarity between two datasets for the overall characteristics preserved in terms of OD links.
  • Trip Length: This metric considers the travel distances of trajectories by using the total length of link trajectories.
  • Radius: In human mobility patterns, the spatial range of daily movements is an important metric. We examined the user’s radius of gyration within the controlled road network area.
  • Gravity: The gravity summary metric examines the impact of the gravity sampling model that was introduced for MobilityGPT at high-level mobility patterns with the same JSD method.
Connectivity: The connectivity metric quantifies the percentage of trajectories that are fully connected from origin to destination, owing to the RCM that samples the next road link conditioned on the connected road links, thereby capturing the underlying road network topology and connectivity patterns.
B.3 Model Details
MobilityGPT utilizes a minimal version of the standard decoder-only transformer architecture (Radford et al., 2019), built upon the minimal GPT implementation from (Karpathy, ). Table 3 summarizes the important model and training parameters employed in this work. Since the Porto and Beijing datasets exhibit different trajectory lengths and characteristics, we utilize distinct block sizes for training and maximum trajectory lengths for synthetic trajectory generation. However, we employ the same set of hyperparameters across both datasets for all training stages, including the reward model and PPO training.
Table 3: Model and training parameters utilized in MobilityGPT
       Parameter Value
N. layers 6
N. heads 4
N embeddings 64
Batch size 64
Learning rate 1e-5
Training size 80%
Test size 20%
Training steps 3000
Block size (Porto) 300
Block size (Beijing) 60
Max. Traj. length (Porto) 278
Max. Traj. length (Beijing) 60
Appendix C Additional Experiments
C.1 Ablation Study on Beijing Dataset
An additional ablation study on the Beijing dataset, shown in Table C.1, investigated the impact of MobilityGPT’s key components. Similar to the Porto dataset, the ablation study on the Beijing dataset exhibits the synergistic effects of MobilityGPT’s novel components – RCM, Gravity Model, and RLTF fine-tuning – in accurately modeling diverse mobility characteristics, boosting performance on key metrics like OD and Trip Length while the connectivity priors ensure accurate capture of network-level patterns.
Table 4: Ablation on Beijing dataset study for different components of MobilityGPT.
       RCM Gravity Model RLTF OD Trip Length Connectivity
0.102 0.201 0.330
0.103 0.151 0.367
0.093 0.145 1.0
0.099 0.131 1.0
0.099 0.123 1.0
C.2 Effect of temperature on MobilityGPT inferencing
The temperature parameter in LLMs, including MobilityGPT, influences the randomness in generating outputs. At lower temperatures, the model behavior is more deterministic, often selecting the most likely next token. As the temperature increases, the model introduces greater randomness, leading to a diverse array of outputs.
Interestingly, as illustrated in Figure 5, MobilityGPT demonstrates a clear trend: its performance tends to increase with higher temperature settings. This observation suggests that mobility data generation can benefit from the randomness of GPT-like models.
Figure 5: Performance of MobilityGPT under different temperature settings.
C.3 Tokenizer Modeling
In the development of the MobilityGPT tokenizer model, we explore the use of specific tokens to mark the start (<BOT>) and end (<EOT>) of trajectories. Our aim is to assess how these tokens affect the model’s performance in human mobility modeling, particularly for sequential trajectory data. Our findings reveal a key insight: using the <EOT> token alone significantly improves performance. This result indicates that emphasizing the end of a trajectory is more beneficial than marking its beginning with a separate token for our mobility modeling task. The effectiveness of the <EOT> token emphasizes the importance of nuanced data representation and the need for task-specific tokenization strategies in sequential data generation with transformer models. An ablation study, detailed in Table 5, compares single and dual-token approaches using various metrics. This study shows that using a single token type outperforms the dual-token method in the MobilityGPT workflow, highlighting the advantages of customized tokenization.
Table 5: Ablation study on the effectness of <EOT>.
       <BOT> <EOT> OD Trip Length Radius Gravity
0.110 0.151 0.128 0.235
0.114 0.124 0.107 0.225
Appendix D Computational Demand
Evaluating computational demand is essential for ensuring the practical feasibility, scalability, and environmental sustainability of generative models. To facilitate a fair comparison, we trained all benchmark models and MobilityGPT on the same computational resources, utilizing an NVIDIA TITAN RTX GPU with 24GB memory, and the results are depicted in Table 6. MobilityGPT balances model size and efficiency: $5.5M$ params, lower than TS-TrajGen ($23.2M$ $7.6M$) but higher than LSTM-TrajGAN ($100k$). Its $\sim$5hr runtime outperforms TS-TrajGen ($>200hr$) and DiffTraj ($48-60hr$) by an order of magnitude, despite LSTM-TrajGAN’s $30min-1hr$ efficiency at the cost of poor predictive performance.
Table 6: Computational performance of different models vs MobilityGPT
  N_param Total running time
Model BJ Porto BJ Porto
SeqGAN 2.6M 800 $\sim$ 12hr $\sim$ 5hr
LSTM-TrajGAN 100k 100k $\sim$ 30mins $\sim$ 1hr
TS-TrajGen 23.2M 7.6M $>$ 200hr $>$ 200hr
DiffTraj 15.7M 15.7M $\sim$ 48hr $\sim$ 60hr
MobilityGPT 5.5M 5.5M $\sim$ 5hr $\sim$ 5hr
MobilityGPT achieves a compelling balance between model capacity and scalability through its moderate size and computational efficiency. Despite having fewer parameters than some benchmarks, MobilityGPT achieves state-of-the-art performance across datasets, requiring substantially less time and resources. Unlike LSTM-TrajGAN, which trades off predictive accuracy for efficiency, MobilityGPT demonstrates the feasibility of combining high performance and practical efficiency within a single model. This makes MobilityGPT a promising choice for practical deployments where both predictive power and computational feasibility are essential.